# Computation of stresses in jammed packings modeled with Tresca friction

### v1

@@ 4 Numerical results
We first describe how the normal stresses are computed in practice. Then, we present a few tests to ensure the consistency of the proposed method. Finally, we show that the method can handle any Voronoi meshes.
@@ 4.2 Implementation
The implementation is performed in Python. The package pyvoro, which uses the library Voro++ [15] is used to generate the Voronoi mesh. The package cvxopt [17] is used to solve (16) and provide the normal stresses by computing the dual linear program. Finally, Firedrake [9] is used to solve (15). The implementation is available in [11].
@@ 4.4 Shear test
Note that the stresses are discontinuous at the interfaces between bricks since one only has $\sigma\in H(\mathrm{div};c)$, for all $c\in\mathcal{M}$. Also, one notices that the stress tensors are not completely symmetric.
@@ 4.5 Homogeneous compression test
We notice that the recomputed stress $\sigma$ in Figure 11 is not homogeneous. Let us consider the constant stress $\sigma=-I_{2}$, since it would seem as a good guess for this test case. Upon recomputing the forces $\bm{f}$ from (14) and $\sigma=-I_{2}$, we notice that $\sigma=-I_{2}$ is also an admissible maximizer of (13) for this test case. It is just not the solution found by the algorithm used in cvxopt. Indeed, this configuration is statically indeterminate and there are several force configurations satisfying the conditions for equilibrium. Note that some stress concentration is observed in Figure 11 since $|\sigma_{ij}|$ can get larger than $|\bm{g}|=1$. Finally, even though the boundary condition is a homogeneous compression, we observe that several cells are in tension. We believe that even though this result sounds counter-intuitive, it is logical since the friction between neighboring cells can lead to having cells under local tension. This is a well documented phenomena. Local tension is required for a crack to open which itself is required for a crack to propagate. In [10], local tension is observed in rocks under compression when they fracture.

### v2

@@ 4 Numerical results
We first describe how the normal stresses are computed in practice. Then, we present a few tests to ensure the consistency of the proposed method. Finally, we show the robustness of the method on unstructured meshes. Note that, even though our methodology only requires convex polygonal cells, the numerical examples below use Voronoi meshes as this is a convenient way to generate structured/unstructured polygonal meshes with convex cells.
@@ 4.2 Implementation
The implementation is performed in Python. The package pyvoro, which uses the library Voro++ [17] is used to generate the Voronoi mesh. The package cvxopt [20] is used to solve (16) and provide the normal stresses by computing the dual linear program. Finally, Firedrake [11] is used to solve (15).
@@ 4.4 Shear test
Note that the stresses are discontinuous at the interfaces between bricks since one only has $\sigma\in H(\mathrm{div};c)$, for all $c\in\mathcal{M}$.
@@ 4.5 Homogeneous compression test
We notice that the recomputed stress $\sigma$ in Figure 11 is not homogeneous.
5 Discussion and Future Work
The numerical results presented in Section 4 highlight several key features of the proposed model which warrant further discussion in light of the physical theory of granular packings.
In the homogeneous compression test (Figure 11), we observed that the solver identifies a non-homogeneous stress state, despite the boundary conditions and domain being consistent with a uniform stress state $\sigma=-I_{2}$. As discussed in Section 2.2, this is a direct consequence of the static indeterminacy (or hyperstaticity) of jammed packings. The uniform solution $\sigma=-I_{2}$ is indeed a valid equilibrium solution (as verified by re-computing forces), but it is just one point in a larger null space of valid states. Our optimization approach selects a specific solution from this manifold, namely a global minimizer for the frictional energy (8). This selection principle mimics nature’s tendency to settle into specific energy-minimized configurations, even if there are other static equilibria. Furthermore, it is well known in optimization theory and statistics that $L^{1}$ minimization promotes sparsity [18, 7]. In the context of granular materials, this mathematical property has a direct physical interpretation: the formation of force chains [13]. By minimizing the sum of frictional work terms (an $L^{1}$-type objective), the system naturally selects a sparse network of load-bearing paths, leaving the majority of the domain in a state of low or zero stress. This aligns with experimental observations of heterogeneous force networks in jammed packings [13].
Regarding torque balance, in the shear test (Figure 9), we noted that $\sigma_{12}\neq\sigma_{21}$. This asymmetry is expected and physically consistent with our modeling choices. In standard limit analysis, one enforces both force and torque balance on rigid blocks. However, in our simplification, we explicitly exclude rotational degrees of freedom. Consequently, our “equilibrium” satisfies weak force balance but not local moment balance. The skew-symmetric component of the reconstructed stress tensor effectively accounts for the residual torques in the system.
Another counter-intuitive result from the compression test is the presence of cells in tension (positive principal stresses), despite the global compressive loading. This is not a numerical artifact but a physical phenomenon observed in frictional granular assemblies. Friction allows for “interlocking” or “bridging” effects where local force chains can shield certain regions or induce local tension perpendicular to the primary load path. This local tension is a critical precursor to fracture initiation in geomaterials [12].
Finally, a natural extension of this work is to restore full mechanical consistency by including rotational degrees of freedom $\theta_{c}$. This would require:
  1. 1.
    Updated Constraints: With rotations, the normal displacement varies linearly along an edge. To ensure non-interpenetration along the entire edge, it is necessary and sufficient to enforce the constraint at both endpoints of the edge. This effectively doubles the number of inequality constraints.
  2. 2.
    Complex Objective Function: Similarly, the tangential relative displacement varies linearly along the edge. The frictional work is the integral of the absolute value of this linear function. While more complex than the constant case, this convex functional can still be linearized (e.g., by splitting the integral at the zero-crossing of the displacement), allowing us to recast as a finite dimensional convex optimization problem.
Implementing these features to recover a fully moment-balanced (symmetric stress) limit analysis remains a challenging avenue for future research. Beyond moment balance, adopting a more realistic friction model such as Coulomb friction is also desirable. While Tresca friction allows for a global energy minimization formulation (since the threshold $s_{T}$ is fixed), Coulomb friction introduces a coupling between the tangential threshold and the normal stress ($|\sigma_{t}|\leq\mu|\sigma_{n}|$) resulting in a Quasi-Variational Inequality [3]. This dependency typically breaks the convexity of the associated energy functional, potentially requiring iterative fixed-point algorithms or incremental evolution schemes rather than a single static optimization step.
Acknowledgements
The authors would like to thank the anonymous referee for their useful comments which helped improve this manuscript. SV was supported in part by NSF grants DMS-2108124 and DMS-2511503.